OpenStudy (anonymous):

You are taking a road trip in a car without A/C. The temperture in the car is 108 degrees F. You buy a cold pop at a gas station. Its initial temperature is 45 degrees F. The pop's temperature reaches 60 degrees F after 41 minutes. Given that T-A/To-A=e^-kt. where T= the temperature of the pop at time t. To= the initial temperature of the pop. A= the temperature in the car. k= a constant that corresponds to the warming rate. and t= the length of time that the pop has been warming up. How many minutes will it take the pop to reach 81 degrees F?

OpenStudy (anonymous):

oh lord really? we can do this but it is going to take some time

OpenStudy (anonymous):

tell me about it. Ive tried it a few already

OpenStudy (anonymous):

ok first off forget about the verbiage

OpenStudy (anonymous):

ok should we set it up first?

OpenStudy (anonymous):

\[108-45=63\] and \[108-60=48\] those are the numbers we will be working with

OpenStudy (anonymous):

we start counting at time t = 0 and at t = 0 we have 63 at t = 41 we have 48 so the problem is find the exponential function containing the points (0, 63) and (41,48)

OpenStudy (anonymous):

so we have \[y=63e^{rt}\] and we are looking for r we know when t = 41 we have y = 48 so write \[48=63e^{41r}\] and solve for r

OpenStudy (anonymous):

solving here is a little tricky. should i divide by 63 then multiply both sides by natural log?

OpenStudy (anonymous):

right i screwed that up! you are right except please don't say to your teacher "multiply by the log" he or she will think you are a moron. say "take the log" and they will be happier. write \[\frac{48}{63}=e^{41r}\] \[\ln(\frac{48}{63})=41r\] \[r=\frac{\ln(\frac{48}{63})}{41}\]

OpenStudy (anonymous):

so now we have r and we can compute it with a calculator

OpenStudy (anonymous):

its -.0066

OpenStudy (anonymous):

-0.006632529

OpenStudy (anonymous):

yeah got it. ok now we want to know when the thing will be 81 degrees but BE CAREFUL because it is the difference in the temperatures we are looking at

OpenStudy (anonymous):

so the difference is 27?

OpenStudy (anonymous):

right. exactly!

OpenStudy (anonymous):

so we want to solve \[63e^{-.0066t}=27\] for t

OpenStudy (anonymous):

and we run through the same routine again

OpenStudy (anonymous):

divide by 63, take the log, divide by -.0066

OpenStudy (radar):

Can you calculate k the warming rate from the fact the pop goes from 45 to 60 (15 degrees) in 41 minutes, or are you not concerned with k??

OpenStudy (anonymous):

@radar no, because it is the difference in the temperatures that decays

OpenStudy (anonymous):

128.3684

OpenStudy (anonymous):

we got k = -.0066 only i called it r . same thing.

OpenStudy (radar):

Oh O.K.

OpenStudy (anonymous):

@bbaker i got the same thing. there is another way to do this that is perhaps quicker but never mind it involves basically the same work

OpenStudy (anonymous):

it is not accepting that answer. I don't know why

OpenStudy (anonymous):

always work with the differences in the temperature and then you don't have to worry so much about the cumbersome formula maybe we made a mistake let me check

OpenStudy (anonymous):

maybe it mean how many MORE minutes will it take, in which case subtract 41

OpenStudy (anonymous):

hmmm. not it either. Im stumped haha

OpenStudy (anonymous):

even using more decimal place accuracy i still get 128 minutes rounded

OpenStudy (anonymous):

ok. I will have to email my teacher. Thank you

OpenStudy (anonymous):

sorry we did not get an answer that the system likes but i don't see a mistake.